# Quasi-Bayes empirical Bayes: a sequential approach to the Poisson compound decision problem

### v1

@@ C Proofs for Section 2
$\liminf_{n}E^{*}\left(\sum_{i=1}^{d}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}\mathbf{h}_{i}(\mathbf{g}_{n})\right)=\liminf_{n}E^{*}\left(\sum_{i=1}^{d}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}\left(\sum_{y\in\mathbb{N}_{0}}\frac{k_{y}(\vartheta_{i})\mathbf{g}_{n,i}p_{\mathbf{g}^{*}}(y)}{p_{\mathbf{g}_{n}}(y)}-\mathbf{g}_{n,i}\right)\right)$
$=\liminf_{n}E^{*}\left(\sum_{y\in\mathbb{N}_{0}}\frac{p_{\mathbf{g}^{*}}(y)^{2}}{p_{\mathbf{g}_{n}}(y)}-1\right)$
$>\liminf_{n}E^{*}\left(\left(\sum_{y\in\mathbb{N}_{0}}\frac{p_{\mathbf{g}_{n}}(y)}{p_{\mathbf{g}^{*}}(y)}p_{\mathbf{g}^{*}_{i}}(y)\right)^{-1}-1\right)=0,$
where the strict inequality holds since, under the contradiction assumption, $\mathds{P}(p_{\mathbf{g}_{n}}\not\rightarrow p_{\mathbf{g}^{*}})>0$. Thus there exists a constant $c>0$ such that
$E^{*}\left(-\sum_{i=1}^{d}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})\right)<-c\alpha_{n+1}$
for $n$ large enough. On the other hand, denoting by $E^{*}_{n}$ the expectation under $\mathds{P}^{*}$, conditional to $\mathcal{G}_{n}$, we can write that
$E^{*}\left(\sum_{i=1}^{d}\frac{\mathbf{g}^{*}_{i}}{(\overline{\mathbf{g}}_{n,i})^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)$
$\leq E^{*}\left(\sum_{i=1}^{d}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)+E^{*}\left(\sum_{i=1}^{d}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n+1,i}^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)$
$\leq\alpha_{n+1}^{2}E^{*}\left(\sum_{i=1}^{d}\mathbf{g}^{*}_{i}\left(\frac{k_{Y_{n+1}}(\vartheta_{i})}{p_{\mathbf{g}_{n}}(Y_{n+1})}-1\right)^{2}\right)+\alpha_{n+1}^{2}E^{*}\left(\sum_{i=1}^{d}\mathbf{g}^{*}_{i}\frac{\left(\frac{k_{Y_{n+1}}(\vartheta_{i})}{p_{\mathbf{g}_{n}}(Y_{n+1})}-1\right)^{2}}{\left(1+\alpha_{n+1}\left(\frac{k_{Y_{n+1}}(\vartheta_{i})}{p_{\mathbf{g}_{n}}(Y_{n+1})}-1\right)\right)^{2}}\right)$
$\leq\alpha_{n+1}^{2}E^{*}\left(1+E^{*}_{n}\left(\sum_{i=1}^{d}\left(\frac{k_{Y_{n+1}}(\vartheta_{i})}{\sum_{i=1}^{d}k_{Y_{n+1}}(\vartheta_{i})\mathbf{g}_{n,i}}\right)^{2}\right)\right)$
@@
$\quad\quad+\frac{\alpha_{n+1}^{2}}{(1-\alpha_{n+1})^{2}}E^{*}\left(1+E^{*}_{n}\left(\sum_{i=1}^{d}\left(\frac{k_{Y_{n+1}}(\vartheta_{i})}{\sum_{i=1}^{d}k_{Y_{n+1}}(\vartheta_{i})\mathbf{g}_{n,i}}\right)^{2}\right)\right)$
@@
$\leq 2\frac{\alpha_{n+1}^{2}}{(1-\alpha_{n+1})^{2}}\left(1+\sum_{y\in\mathbb{N}_{0}}\sum_{i,j,k}\frac{k_{y}(\vartheta_{i})^{2}}{(\sum_{\vartheta_{k}}k_{y}(\vartheta_{k})g_{n}(\vartheta_{k}))^{2}}k_{y}(\vartheta_{j})g^{*}(\vartheta_{j})\right)$
$\leq 2\frac{\alpha_{n+1}^{2}}{(1-\alpha_{n+1})^{2}}\left(1+\sum_{y\in\mathbb{N}_{0}}\sum_{i,j,k}\frac{k_{y}(\vartheta_{i})^{2}}{k_{y}(\vartheta_{k})^{2}}k_{y}(\vartheta_{j})g^{*}(\vartheta_{j})g_{n}(\vartheta_{k})\right)$
$\leq 2\frac{\alpha_{n+1}^{2}}{(1-\alpha_{n+1})^{2}}\left(1+d\sum_{y\in\mathbb{N}_{0}}\ max_{i,j,k}\frac{k_{y}(\vartheta_{i})^{2}}{k_{y}(\vartheta_{k})^{2}}k_{y}(\vartheta_{j})\right)$
$=o(\alpha_{n}),$
@@
where the last equality holds since
$\sum_{y\in\mathbb{N}_{0}}\frac{k_{y}(\vartheta_{i})^{2}}{k_{y}(\vartheta_{j})^{2}}k_{y}(\vartheta_{k})=\sum_{y\in\mathbb{N}_{0}}e^{2\vartheta_{j}-2\vartheta_{i}-\vartheta_{k}}\left(\frac{\vartheta_{i}^{2}\vartheta_{k}}{\vartheta_{j}^{2}}\right)^{y}\frac{1}{y!}<\infty.$
It follows that,
$E^{*}\left(\mathcal{L}(\mathbf{g}_{n+1})\right)\leq E^{*}\left(\mathcal{L}(\mathbf{g}_{n})\right),$
@@
for $n$ large enough, which contradicts $E(\mathcal{L}(\mathbf{g}_{n}))\rightarrow\infty$. ∎
@@
Let $\mathbf{g}_{n}$ and $\mathbf{g}^{*}$ be the representations of $g_{n}$ ($n\geq 0$) and $g^{*}$ in $\mathbb{R}^{d}$. Under the assumptions of Theorem 2.4, for every $\delta<2-1/\gamma$
$||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2}=o(n^{-\delta})\quad\mathds{P}^{*}-a.s.$
@@
In this proof, $\mathbf{g}_{n}$ and $\mathbf{g}^{*}$ are considered as random vectors in $\{\mathbf{v}\in\mathbb{R}^{d}:||\mathbf{v}||\leq 1\}$. By Lemma F.4, $\mathbf{g}_{n}$, seen as vectors in $\Delta^{d}$ converge $P^{*}$-a.s. to $\mathbf{g}\in\Delta^{d}$. This implies that $\mathbf{g}_{n}$ converges to $\mathbf{g}^{*}$ also when seen as vectors in $\mathbb{R}^{d}$. Moreover, $(\mathbf{g}_{n})$ satisfies the stochastic approximation in $\mathbb{R}^{d}$:
@@
$\mathbf{h}(\mathbf{g})=\mathbf{g}\circ\left(\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}}{p_{\mathbf{g}}(y)}p_{\mathbf{g}^{*}}(y)-\bf 1\right),$
@@
${\bm{\epsilon}}_{n+1}=\mathbf{g}\circ\left(\frac{\mathbf{k}_{Y_{n+1}}}{p_{\mathbf{g}}(Y_{n+1})}p_{\mathbf{g}^{*}}(Y_{n+1})-\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}}{p_{\mathbf{g}}(y)}p_{\mathbf{g}^{*}}(y)\right).$
Based on Lemma C.4, $\mathbf{g}_{n}$ converges to $\mathbf{g}^{*}$, $\mathds{P}^{*}$-a.s. The thesis follows from Theorem 3.1.1 in Chen (2005) if we can prove that the following conditions hold:
  • B1
@@
  • B2
    $\sum_{n=1}^{\infty}\alpha_{n}^{1-\delta/2}\epsilon_{n}$ converges, $\mathds{P}^{*}$-a.s.;
  • B3
    $\mathbf{h}$ is measurable and locally bounded, and is differentiable at $\mathbf{g}^{*}$. Let $\mathbf{H}$ be a matrix such that, as $\mathbf{g}\rightarrow\mathbf{g}^{*}$,
    $\mathbf{h}(\mathbf{g})=\mathbf{H}(\mathbf{g}-\mathbf{g}^{*})+\mathbf{r}(g),\quad\mathbf{r}(g^{*})=\mathbf{0},\quad||\mathbf{r}(\mathbf{g})||=o(||\mathbf{g}-\mathbf{g}^{*}||).$
@@
Theorem 3.1.1 in Chen (2005) also requests the following condition: B4 There exists a continuously differentiable function $\mathscr{l}:\Delta^{d}\rightarrow\mathbb{R}$ such that $\sup_{\delta_{1}\leq{\rm d}(\mathbf{g},\mathbf{g}^{*})\leq\delta_{2}}\mathbf{h}(\mathbf{g})^{T}\nabla_{\mathbf{g}}\mathscr{l}<0$ for every $\delta_{2}>\delta_{1}>0$. However, a close inspection of the proof of Theorem 3.1.1 in Chen shows that B4 is employed only to ensure convergence of $\mathbf{g}_{n}$ towards $\mathbf{g}^{*}$. In our case, this property has been proved in Lemma C.4.
B1 is obvious. To prove B2, notice that $(\sum_{k=1}^{n}\alpha_{k}^{1-\delta/2}\epsilon_{k})_{n\geq 1}$ is a martingale. Moreover, there exists a constant $C$ such that, for every $i=1,\dots,d$,
$\sup_{n}E^{*}((\sum_{k=1}^{n}\alpha_{k}^{1-\delta/2}{\bm{\epsilon}}_{k,i})^{2})$ $=\sup_{n}\sum_{k=1}^{n}\alpha_{k}^{2-\delta}E^{*}({\bm{\epsilon}}_{k,i})^{2})$
@@
since $\gamma(2-\delta)>1$ and $({\bm{\epsilon}}_{n,i})$ are uniformly bounded. We now prove B3. The function $\mathbf{h}$ is continuously differentiable at $\mathbf{g}^{*}$ and
$\mathbf{H}_{i,j}=-\mathbf{g}^{*}_{i}\sum_{y\in\mathbb{N}_{0}}\frac{k_{y}(\vartheta_{i})k_{y}(\vartheta_{j})}{p_{g^{*}}(y)}.$
The matrix $\mathbf{H}$ can be seen as the product of the diagonal matrix ${\rm diag}(\mathbf{g}^{*})$, which is positive definite, times $-\mathbf{M}$, with $\mathbf{M}_{i,j}=\sum_{y\in\mathbb{N}_{0}}\frac{k_{y}(\vartheta_{i})k_{y}(\vartheta_{j})}{p_{g^{*}}(y)}$. We now prove that $\mathbf{M}$ is positive definite. Given a vector $\mathbf{u}\in\mathbb{R}^{d}$ with $||\mathbf{u}||=1$, we can write that
$\mathbf{u}^{T}\mathbf{M}\mathbf{u}=$ $\sum_{i,j=1}^{d}\mathbf{u}_{i}\mathbf{u}_{j}\sum_{y\in\mathbb{N}_{0}}\frac{k_{y}(\vartheta_{i})k_{y}(\vartheta_{j})}{p_{g^{*}}(y)}=\sum_{y\in\mathbb{N}_{0}}\frac{1}{p_{g^{*}}(y)}\left(\sum_{i=1}^{d}\mathbf{u}_{i}k_{y}(\vartheta_{i})\right)^{2}\geq 0,$
with equality if and only if $\sum_{i=1}^{d}\mathbf{u}_{i}k_{y}(\vartheta_{i})=0$ for every $y\in\mathbb{N}_{0}$. The only solution to the last equation in $\mathbf{u}$ is the zero vector, which is inconsistent with $||\mathbf{u}||=1$. Since the function $\mathbf{u}^{T}\mathbf{M}\mathbf{u}$ is continuous on the compact set $||\mathbf{u}||=1$, then the minimum is attained, and by the above reasoning it is strictly positive. It follows that $\mathbf{M}$ is positive definite. ∎
@@
Denoting by $(\mathbf{g}_{n})$ and $\mathbf{g}^{*}$ the representations of $(g_{n})$ and $g^{*}$ in $\mathbb{R}^{d}$, we can write, by Lemma C.3 and Lemma C.5,

### v2

@@ C Proofs for Section 2
$\liminf_{n}E_{G^{*}}\left(\sum_{i=1}^{d_{\eta}}\frac{\mathbf{g}_{\eta,i}^{*}}{\mathbf{g}_{n,i}}\mathbf{h}_{i}(\mathbf{g}_{n})\mid\mathcal{G}_{n}\right)=\liminf_{n}\sum_{i=1}^{d_{\eta}}\frac{\mathbf{g}_{\eta,i}^{*}}{\mathbf{g}_{n,i}}\left(\sum_{y\in\mathbb{N}_{0}}\frac{k_{y}(\vartheta_{i})\mathbf{g}_{n,i}p_{G^{*}}(y)}{p_{\mathbf{g}_{n}}(y)}-\mathbf{g}_{n,i}\right)$
$=\liminf_{n}\sum_{i=1}^{d_{\eta}}\mathbf{g}_{\eta,i}^{*}\left(\sum_{y\in\mathbb{N}_{0}}\frac{k_{y}(\vartheta_{i})}{p_{\mathbf{g}_{n}}(y)}p_{G^{*}}(y)-1\right)$
$>0,$
by (C.15). Thus there exists a constant $c>0$ such that
$E_{G^{*}}\left(-\sum_{i=1}^{d_{\eta}}\frac{\mathbf{g}_{\eta,i}^{*}}{\mathbf{g}_{n,i}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})\mid\mathcal{G}_{n}\right)<-c\alpha_{n+1}$
for $n$ large enough. On the other hand,
$E_{G^{*}}\left(\sum_{i=1}^{d_{\eta}}\frac{\mathbf{g}_{\eta,i}^{*}}{(\overline{\mathbf{g}}_{n,i})^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)$
$\leq E_{G^{*}}\left(\sum_{i=1}^{d_{\eta}}\frac{\mathbf{g}_{\eta,i}^{*}}{\mathbf{g}_{n,i}^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)+E_{G^{*}}\left(\sum_{i=1}^{d_{\eta}}\frac{\mathbf{g}_{\eta,i}^{*}}{\mathbf{g}_{n+1,i}^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)$
$\leq\alpha_{n+1}^{2}E_{G^{*}}\left(\sum_{i=1}^{d_{\eta}}\mathbf{g}_{\eta,i}^{*}\left(\frac{k_{Y_{n+1}}(\vartheta_{i})}{p_{\mathbf{g}_{n}}(Y_{n+1})}-1\right)^{2}\right)+\alpha_{n+1}^{2}E_{G^{*}}\left(\sum_{i=1}^{d_{\eta}}\mathbf{g}_{\eta,i}^{*}\frac{\left(\frac{k_{Y_{n+1}}(\vartheta_{i})}{p_{\mathbf{g}_{n}}(Y_{n+1})}-1\right)^{2}}{\left(1+\alpha_{n+1}\left(\frac{k_{Y_{n+1}}(\vartheta_{i})}{p_{\mathbf{g}_{n}}(Y_{n+1})}-1\right)\right)^{2}}\right)$
$\leq\alpha_{n+1}^{2}E_{G^{*}}\left(1+E_{G^{*}}\left(\sum_{i=1}^{d_{\eta}}\left(\frac{k_{Y_{n+1}}(\vartheta_{i})}{\sum_{i=1}^{d_{\eta}}k_{Y_{n+1}}(\vartheta_{i})\mathbf{g}_{n,i}}\right)^{2}\mid\mathcal{G}_{n}\right)\right)$
@@
$\quad\quad+\frac{\alpha_{n+1}^{2}}{(1-\alpha_{n+1})^{2}}E_{G^{*}}\left(1+E_{G^{*}}\left(\sum_{i=1}^{d_{\eta}}\left(\frac{k_{Y_{n+1}}(\vartheta_{i})}{\sum_{i=1}^{d_{\eta}}k_{Y_{n+1}}(\vartheta_{i})\mathbf{g}_{n,i}}\right)^{2}\mid\mathcal{G}_{n}\right)\right)$
@@
$\leq 2\frac{\alpha_{n+1}^{2}}{(1-\alpha_{n+1})^{2}}\left(1+\sum_{y\in\mathbb{N}_{0}}\sum_{i,k}\frac{k_{y}(\vartheta_{i})^{2}}{(\sum_{\vartheta_{k}}k_{y}(\vartheta_{k})g_{n}(\vartheta_{k}))^{2}}\int_{\Theta}k_{y}(\theta)G^{*}(\mathrm{d}\theta)\right)$
$\leq 2\frac{\alpha_{n+1}^{2}}{(1-\alpha_{n+1})^{2}}\left(1+\sum_{i=1}^{d_{\eta}}\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{i})^{2}}{p_{g_{n}}(y)^{2}}p_{G^{*}}(y)\right)$
$=o(\alpha_{n}).$
@@
where the last equality holds by (C.6), as proved in Lemma C.3. It follows that,
$E_{G^{*}}\left(\mathcal{L}(\mathbf{g}_{n+1})\right)\leq E_{G^{*}}\left(\mathcal{L}(\mathbf{g}_{n})\right),$
@@
for $n$ large enough, which contradicts $E_{G^{*}}(\mathcal{L}(\mathbf{g}_{n}))\rightarrow\infty$. ∎
@@
Let $\mathbf{g}_{n}$ and $\mathbf{g}_{\eta}^{*}$ be the representations of $g_{n}$ ($n\geq 0$) and $g_{\eta}^{*}$ in $\mathbb{R}^{d_{\eta}}$. Under the assumptions of Theorem 2.4 with $\gamma<1$, for every $\delta<2-1/\gamma$
$||\mathbf{g}_{n}-\mathbf{g}_{\eta}^{*}||^{2}=o(n^{-\delta})\quad\mathds{P}_{G^{*}}-a.s.$
@@
The random vectors $\mathbf{g}_{n}$ and $\mathbf{g}_{\eta}^{*}$ take values in the compact set $\{\mathbf{v}\in\mathbb{R}^{d_{\eta}}:||\mathbf{v}||\leq 1\}$. By Lemma C.4, $\mathbf{g}_{n}$, seen as vectors in $\Delta^{d_{\eta}}$ converge $P_{G^{*}}$-a.s. to $\mathbf{g}_{\eta}^{*}\in\Delta^{d_{\eta}}$. This implies that $\mathbf{g}_{n}$ converges to $\mathbf{g}_{\eta}^{*}$ also when seen as vectors in $\mathbb{R}^{d_{\eta}}$. Moreover, $(\mathbf{g}_{n})$ satisfies the stochastic approximation in $\mathbb{R}^{d_{\eta}}$:
@@
$\mathbf{h}(\mathbf{g})=\mathbf{g}\circ\left(\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}}{p_{\mathbf{g}}(y)}p_{G^{*}}(y)-\bf 1\right),$
@@
${\bm{\epsilon}}_{n+1}=\mathbf{g}\circ\left(\frac{\mathbf{k}_{Y_{n+1}}}{p_{\mathbf{g}}(Y_{n+1})}-\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}}{p_{\mathbf{g}}(y)}p_{G^{*}}(y)\right).$
The thesis follows from Theorem 3.1.1 in Chen (2005) if we can prove that the following conditions hold:
  • T1
@@
  • T2
    $\sum_{n=1}^{\infty}\alpha_{n}^{1-\delta/2}\epsilon_{n}$ converges, $\mathds{P}_{G^{*}}$-a.s.;
  • T3
    $\mathbf{h}$ is measurable and locally bounded, and is differentiable at $\mathbf{g}_{\eta}^{*}$. Let $\mathbf{H}$ be a matrix such that, as $\mathbf{g}\rightarrow\mathbf{g}_{\eta}^{*}$,
    $\mathbf{h}(\mathbf{g})=\mathbf{H}(\mathbf{g}-\mathbf{g}_{\eta}^{*})+\mathbf{r}(g),\quad\mathbf{r}(g_{\eta}^{*})=\mathbf{0},\quad||\mathbf{r}(\mathbf{g})||=o(||\mathbf{g}-\mathbf{g}_{\eta}^{*}||).$
@@
Theorem 3.1.1 in Chen (2005) also requests the following condition: T4 There exists a continuously differentiable function $\mathscr{l}:\Delta^{d_{\eta}}\rightarrow\mathbb{R}$ such that $\sup_{\delta_{1}\leq{\rm d}(\mathbf{g},\mathbf{g}_{\eta}^{*})\leq\delta_{2}}\mathbf{h}(\mathbf{g})^{T}\nabla_{\mathbf{g}}\mathscr{l}<0$ for every $\delta_{2}>\delta_{1}>0$. However, a close inspection of the proof of Theorem 3.1.1 in Chen shows that T4 is employed only to ensure convergence of $\mathbf{g}_{n}$ towards $\mathbf{g}_{\eta}^{*}$. In our case, this property has been proved in Lemma C.4.
T1 is obvious. To prove T2, notice that $(\sum_{k=1}^{n}\alpha_{k}^{1-\delta/2}\epsilon_{k})_{n\geq 1}$ is a martingale. Moreover, there exists a constant $C$ such that, for every $i=1,\dots,{d_{\eta}}$,
$\sup_{n}E_{G^{*}}((\sum_{k=1}^{n}\alpha_{k}^{1-\delta/2}{\bm{\epsilon}}_{k,i})^{2})$ $=\sup_{n}\sum_{k=1}^{n}\alpha_{k}^{2-\delta}E_{G^{*}}({\bm{\epsilon}}_{k,i})^{2})$
@@
since $\gamma(2-\delta)>1$ and $({\bm{\epsilon}}_{n,i})$ are uniformly bounded. We now prove T3. The function $\mathbf{h}$ is continuously differentiable at $\mathbf{g}_{\eta}^{*}$ and
$\mathbf{H}_{i,j}=-\mathbf{g}_{\eta,i}^{*}\sum_{y\in\mathbb{N}_{0}}\frac{k_{y}(\vartheta_{i})k_{y}(\vartheta_{j})}{p_{g_{\eta}^{*}}(y)^{2}}p_{G^{*}}(y).$
The matrix $\mathbf{H}$ can be seen as the product of the diagonal matrix ${\rm diag}(\mathbf{g}_{\eta}^{*})$, which is positive definite, times $-\mathbf{M}$, with $\mathbf{M}_{i,j}=\sum_{y\in\mathbb{N}_{0}}\frac{k_{y}(\vartheta_{i})k_{y}(\vartheta_{j})}{p_{g_{\eta}^{*}}(y)^{2}}p_{G^{*}}(y)$. We now prove that $\mathbf{M}$ is positive definite. Given a vector $\mathbf{u}\in\mathbb{R}^{d_{\eta}}$ with $||\mathbf{u}||=1$, we can write that
$\mathbf{u}^{T}\mathbf{M}\mathbf{u}=$ $\sum_{i,j=1}^{d_{\eta}}\mathbf{u}_{i}\mathbf{u}_{j}\sum_{y\in\mathbb{N}_{0}}\frac{k_{y}(\vartheta_{i})k_{y}(\vartheta_{j})}{p_{g_{\eta}^{*}}(y)^{2}}p_{G^{*}}(y)=\sum_{y\in\mathbb{N}_{0}}\frac{p_{G^{*}{(y)}}}{p_{g_{\eta}^{*}}(y)^{2}}\left(\sum_{i=1}^{d_{\eta}}\mathbf{u}_{i}k_{y}(\vartheta_{i})\right)^{2}\geq 0,$
with equality if and only if $\sum_{i=1}^{d_{\eta}}\mathbf{u}_{i}k_{y}(\vartheta_{i})=0$ for every $y\in\mathbb{N}_{0}$. The only solution to the last equation in $\mathbf{u}$ is the zero vector, which is inconsistent with $||\mathbf{u}||=1$. Since the function $\mathbf{u}^{T}\mathbf{M}\mathbf{u}$ is continuous on the compact set $||\mathbf{u}||=1$, then the minimum is attained, and by the above reasoning it is strictly positive. It follows that $\mathbf{M}$ is positive definite. ∎
Lemma C.7.
${\rm Regret}(g_{n},g_{\eta}^{*})\leq C||\mathbf{g}_{n}-\mathbf{g}_{\eta}^{*}||^{2},$
where $\mathbf{g}_{n}$ and $\mathbf{g}_{\eta}^{*}$ are the representations of $g_{n}$ and $g_{\eta}^{*}$ in $\mathbb{R}^{d_{\eta}}$.
Proof.
We can write that
$\sum_{y\in\mathbb{N}_{0}}(\hat{\theta}_{g_{n}}(y)-\hat{\theta}_{g_{\eta}^{*}}(y))^{2}p_{g_{\eta}^{*}}(y)$
$=\sum_{y\in\mathbb{N}_{0}}\left(\frac{\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y+1}e^{-\vartheta_{i}}g_{n}(\vartheta_{i})}{\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y}e^{-\vartheta_{i}}g_{n}(\vartheta_{i})}-\frac{\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y+1}e^{-\vartheta_{i}}g_{\eta}^{*}(\vartheta_{i})}{\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y}e^{-\vartheta_{i}}g_{\eta}^{*}(\vartheta_{i})}\right)^{2}p_{G^{*}}(y)$
$\leq\sum_{y\in\mathbb{N}_{0}}\frac{\left(\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y+1}e^{-\vartheta_{i}}\left(g_{n}(\vartheta_{i})-g_{\eta}^{*}(\vartheta_{i})\right)\right)^{2}}{(\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y}e^{-\vartheta_{i}}g_{\eta}^{*}(\vartheta_{i}))^{2}}p_{G^{*}}(y)$
$+\sum_{y\in\mathbb{N}_{0}}\frac{\left(\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y+1}e^{-\vartheta_{i}}g_{n}(\vartheta_{i})\right)^{2}\left(\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y}e^{-\vartheta_{i}}(g_{n}(\vartheta_{i})-g_{\eta}^{*}(\vartheta_{i})\right)^{2}}{\left(\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y}e^{-\vartheta_{i}}g_{n}(\vartheta_{i})\right)^{2}\left(\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y}e^{-\vartheta_{i}}g_{\eta}^{*}(\vartheta_{i})\right)^{2}}p_{G^{*}}(y)$
$\leq 2\sum_{y\in\mathbb{N}_{0}}\frac{\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{2y+2}e^{-2\vartheta_{i}}||\mathbf{g}_{n}-\mathbf{g}_{\eta}^{*}||^{2}}{\left(\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y}e^{-\vartheta_{i}}g_{\eta}^{*}(\vartheta_{i})\right)^{2}}p_{G^{*}}(y)+2\sum_{y\in\mathbb{N}_{0}}\frac{\vartheta_{d_{\eta}}^{2}\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{2y}e^{-2\vartheta_{i}}||\mathbf{g}_{n}-\mathbf{g}_{\eta}^{*}||^{2}}{\left(\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y}e^{-\vartheta_{i}}g_{\eta}^{*}(\vartheta_{i})\right)^{2}}p_{G^{*}}(y)$
$\leq 4{d_{\eta}}\vartheta_{d_{\eta}}^{2}e^{2(\vartheta_{d_{\eta}}-\vartheta_{1})}||\mathbf{g}_{n}-\mathbf{g}_{\eta}^{*}||^{2}\int_{0}^{\infty}e^{-\theta}\sum_{y\in\mathbb{N}_{0}}\frac{1}{y!}\left(\frac{\theta\vartheta_{d_{\eta}}^{2}}{\vartheta_{1}^{2}}\right)^{y}G^{*}(d\theta)$
$\leq 4{d_{\eta}}\vartheta_{d_{\eta}}^{2}e^{2(\vartheta_{d_{\eta}}-\vartheta_{1})}||\mathbf{g}_{n}-\mathbf{g}_{\eta}^{*}||^{2}\int_{0}^{\infty}e^{\theta({d_{\eta}}^{2}-1)}G^{*}(d\theta)$
$\leq C||\mathbf{g}_{n}-\mathbf{g}_{\eta}^{*}||^{2},$
with $C=4{d_{\eta}}\vartheta_{d_{\eta}}^{2}e^{2(\vartheta_{d_{\eta}}-\vartheta_{1})}\int_{0}^{\infty}e^{\theta({d_{\eta}}^{2}-1)}G^{*}(d\theta)<+\infty,$ by (C.8). Thus,
${\rm Regret}(g_{n},g_{\eta}^{*})\leq C||\mathbf{g}_{n}-\mathbf{g}_{\eta}^{*}||^{2},$
with $C<+\infty$. ∎
@@
Let $\mathbf{g}_{n}$ and $\mathbf{g}_{\eta}^{*}$ denote the representation of $g_{n}$ and $g_{\eta}^{*}$ in $\Delta_{\eta}=\{\mathbf{v}=(\mathbf{v}_{1},\dots,\mathbf{v}_{d_{\eta}-1}):\mathbf{v}_{i}\geq 0,\sum_{i=1}^{d_{\eta}-1}\mathbf{v}_{i}\leq 1\}$. The sequence $(\mathbf{g}_{n})$ satisfies the stochastic approximation (C.13), with $\mathbf{g}_{0}\in\mathring{(\Delta_{\eta})}$. By Lemma C.4, $\mathbf{g}_{n}$ converges to $\mathbf{g}_{\eta}^{*}$, $\mathds{P}^{*}$-a.s. It follows that $g_{n}(\vartheta_{j})$ converges to $g_{\eta}^{*}(\vartheta_{j})$ $\mathds{P}^{*}$-a.s. for every $\vartheta_{j}\in\Theta_{\eta}$, which imples that $\hat{\theta}_{g_{n}}(y)$ converges to $\hat{\theta}_{g^{*}_{\eta}}(y)$ for every $y\in\mathbb{N}_{0}$.
Now assume that $\gamma<1$. Denoting by $\mathbf{g}_{n}$ and $\mathbf{g}_{\eta}^{*}$ the representations of $g_{n}$ and $g_{\eta}^{*}$ in $\mathbb{R}^{d_{\eta}}$, we can write, by Lemma C.5 that
$||\mathbf{g}_{n}-\mathbf{g}_{\eta}^{*}||^{2}=o(n^{-\delta}),$
for every $\delta<2-1/\gamma$. By Lemma C.7,
${\rm Regret}(g_{n},g_{\eta}^{*})\leq C||\mathbf{g}_{n}-\mathbf{g}_{\eta}^{*}||^{2}=o(n^{-\delta}).$
$\square$